# thanks: Sample text

### v1

@@ III-C Certification
Stationarity $S_{\lambda}Y=\bf{0}$ $S_{\lambda}Z=\bf{0}$
@@
Dual feasibility not relevant $S_{\lambda}\succeq 0$
TABLE I: Optimality conditions for the BM and SDP problems, where $S_{\lambda}\triangleq Q+\sum\lambda_{i}A_{i}$. The conditions for BM are simply necessary for first-order optimality, while the conditions for the SDP are necessary and sufficient for global optimality. Dual feasibility in BM is not relevant, as it is not needed to construct the certification scheme we describe. Both sets of conditions can be found from the KKT conditions of the respective problems [2].
@@
As described in Table I, comparison of the KKT conditions of the BM and SDP problems reveals that a first-order stationary point of the BM problem $Y^{*}$ can generate a candidate solution to the SDP problem $Z^{*}=Y^{*}(Y^{*})^{\top}$, which automatically satisfies all SDP optimality conditions except for the dual feasibility condition $S_{\lambda}\succeq 0$. As a result, a BM solution can be certified as globally optimal by checking positive semidefiniteness of the certificate matrix,
@@
where $Q$ is the data matrix describing the original QCQP, and $\lambda_{i}$ and $A_{i}$ are the Lagrange multipliers and constraint matrices. In practice, the fastest and most reliable way to evaluate $S_{\lambda}\succeq 0$ is to compute a Cholesky factorization of $S_{\lambda}+\epsilon I$ for a small $\epsilon>0$, which will fail if $S_{\lambda}$ is not positive semidefinite.222 If $S_{\lambda}$ is not positive semidefinite, then a negative eigenpair of $S_{\lambda}$ can be used to construct a second-order descent direction to kickstart optimization at the next level ($r+1$) of the BM hierarchy. See e.g., [10] for more details.
@@ IV Applications
Geometric structure. Perception problems typically possess smooth, geometric structure that is naturally expressed as well-studied manifolds. For example, orthogonality and unit-norm constraints appear throughout perception – correspondingly the Stiefel manifold is ubiquitous in certifiable perception. In fact, all existing works using BM for certifiable perception can be expressed as optimization over the Stiefel manifold and Euclidean space.333 Note that the unit-sphere is a special case of the Stiefel manifold. Karimian and Tron [34] are a notable partial exception, as they derive a custom manifold by adding an additional explicit constraint to the Stiefel manifold.

### v2

@@ III-C Certification
Stationarity $S_{\lambda}Y=\mathbf{0}$ $S_{\lambda}=Q+\sum\lambda_{i}A_{i}$
Complementarity $S_{\lambda}Z=\mathbf{0}$
@@
Dual feasibility $S_{\lambda}\succeq 0$
TABLE I: Optimality conditions for the BM and SDP problems, where $S_{\lambda}\triangleq Q+\sum\lambda_{i}A_{i}$. A “—” in an entry indicates that the condition is not relevant to the theory presented in this paper. The BM conditions are necessary for first-order optimality, while the SDP conditions are necessary and sufficient for global optimality. All conditions can be derived from the first-order optimality conditions of the respective problem [2]. When all SDP conditions are satisfied, $Z$ and $S_{\lambda}$ are a primal-dual pair of solutions to the SDP relaxation.
@@
As described in Table I, comparison of the KKT conditions of the BM and SDP problems reveals that a first-order stationary point of the BM problem $Y^{*}$ can generate a candidate solution to the SDP problem $Z^{*}=Y^{*}(Y^{*})^{\top}$, which automatically satisfies all SDP optimality conditions except for the dual feasibility condition $S_{\lambda}\succeq 0$. As a result, a BM solution can be certified as globally optimal by checking positive semidefiniteness of the certificate matrix, [^2] Note that when this certificate matrix is positive semidefinite (and thus all the SDP optimality conditions in Table I are satisfied), the certificate matrix is also a solution the SDP dual problem. As a result, the certification scheme described here can also be viewed as searching for a dual solution.
@@
where $Q$ is the data matrix describing the original QCQP, and $\lambda_{i}$ and $A_{i}$ are the Lagrange multipliers and constraint matrices. In practice, the fastest and most reliable way to evaluate $S_{\lambda}\succeq 0$ is to compute a Cholesky factorization of $S_{\lambda}+\epsilon I$ for a small $\epsilon>0$, which will fail if $S_{\lambda}$ is not positive semidefinite.333 If $S_{\lambda}$ is not positive semidefinite, then a negative eigenpair of $S_{\lambda}$ can be used to construct a second-order descent direction to kickstart optimization at the next level ($r+1$) of the BM hierarchy. See e.g., [10] for more details.
@@ IV Applications
Geometric structure. Perception problems typically possess smooth, geometric structure that is naturally expressed as well-studied manifolds. For example, orthogonality and unit-norm constraints appear throughout perception – correspondingly the Stiefel manifold is ubiquitous in certifiable perception. In fact, all existing works using BM for certifiable perception can be expressed as optimization over the Stiefel manifold and Euclidean space.444 Note that the unit-sphere is a special case of the Stiefel manifold. Karimian and Tron [34] are a notable partial exception, as they derive a custom manifold by adding an additional explicit constraint to the Stiefel manifold.